OpenStudy (anonymous):

The speed of sound V near the Earth's surface can be represented by this equation,v=20√t+273 , where t is the temperature in degrees Celsius. Find the temperature when V = 380 meters per second. Round to the nearest whole number. A. 88 B. 90 C. 92 D. 94

OpenStudy (dumbcow):

none of those are right...it should be close to 30

OpenStudy (anonymous):

\[380=20\sqrt{t+273} \implies 19^2=t+273 \implies t=88\]

OpenStudy (anonymous):

The notation of the problem statement wasn't exactly as clear as it might have been, but amazing as it might seem, I remembered that formula from an aeronautical engineering class about a lifetime ago.